find the partial decomposition for the rational expression. 3/2x(3x+1) 3/2x(3x+1)=?/2x+?/3x+1 the question mark represent a blank space
Answer by stanbon(57387)

find the partial decomposition for the rational expression. 3/2x(3x+1) 3/2x(3x+1)=?/2x+?/3x+1 the question mark represent a blank space

------------
3/2x(3x+1) = A/2x + B/(3x+1)
= [A(3x+1)+B(2x)]/[2x(3x+1)]
Since the denominators are equal the numerators must be equal:
[3Ax+A + 2Bx] = 3
(3A+2B)x + A = 0x + 3
A = 3
Therefore:
3A+2B = 0
3*3 + 2B = 0
2B = - 9
B = -4.5
===================
Final Answer:
3/2x(3x+1) = 3/2x - 4.5/(3x+1)
================
Cheers,
Stan H.